OpenStudy (anonymous):

Choose the statement that correctly lists equivalent ratios. A. 1 : 6 = 20 : 120 = 5 : 30 B. 1 : 6 = 4 : 18 = 6 : 36 C. 1 : 8 = 20 : 120 = 3 : 24 D. 1 : 8 = 3 : 32 = 5 : 48

OpenStudy (cwrw238):

hint: these are correct 1 : 2 = 8 : 16 = 4 : 8 because the second one is obtained by multiplying the 1 and 2 by 8 and the third by multiplying 1 and 2 by 4 or you can test it by dividing: 1 / 2 = 8/16 = 4 / 8

OpenStudy (anonymous):

Thanks !

OpenStudy (cwrw238):

yw